How do i solve the question and whats the answer for it. I'd like the steps so that i can know in the future how to do similar ones.
x^2 - 2x - 4 = 11

Answer by stanbon(75887)

How do i solve the question and whats the answer for it. I'd like the steps so that i can know in the future how to do similar ones.
x^2 - 2x - 4 = 11
---
Rearrange:
x^2 - 2x - 15 = 0
---
Factor:
(x-5)(x+3) = 0
----
x = 5 or x = -3
=====================
Note: Your problem was factorable. If your problem is not
factorable use the Quadratic Formula to find the solutions.
==============================
Cheers,
Stan H.
=====================